OpenStudy (anonymous):

Graph the quadratic variation if g(x) varies directly with x2, and g(x) = 75 when x = 5.

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

Varies directly means they are proportional.

OpenStudy (anonymous):

So we can say: \[ \frac{g(x)}{g(x_0)} = \frac{x^2}{x_0^2} \implies g(x) = x^2\frac{g(x_0)}{x_0^2} \]

OpenStudy (anonymous):

Is this a statistics question?